evaluate the following functions:
f(x)=-2x^2+3x+x-10
f(1)=? (I got -4)
f(-2)=? (I got -26)
f(3)=? (I got 20)
I just want to make sure I got the right answers and if I didn't where I went wrong. Thanks!

Answer by ice_water35(8)

>>>>>Your correct in the first and second question. But in F(3), there's something wrong.
f(x)=-2x^2+3x+x-10
f(3)=-2(3)^2+3(3)+(3)-10
f(3)=-2(9)+3(3)+3-10
f(3)=-18+9+3-10
f(3)=-9+3-10
f(3)=-6-10
f(3)=-16
if you have further questions, you may contact me. This is my number: 09223011043